media%2F20e%2F20e748ab-388e-4e9f-ba16-50
Answer #1

a) for path C:

total distance travelled = ( 10-2)+ (10-8)+( 11-8) = 8+2+3= 13 m

b) distance from start to finish = (11-2) = 9m

c) displacement from start to finish = (11-2) = 9m